a first order parallel reaction given as"a-----b and c .given that k1 is rate constant for formation of b and k2 is rate constant for c.k1=1.3 * 10 to power -5 ,sec to power -1.also k1\k2=1\9.find conc.ratio of c to a if an experiment is allowed to start with only a for 1 hr. 

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi,

A---->B     k1

A---->C     k2

d[B]/dt = k1*[A]      d[B] = k1*[A] dt

d[C]/dt = k2*[A]      d[C] = k2*[A] dt

therefore [B]/[C] = k1/k2 = 1/9

[A] = [B]+[C]

[A]/[C] = [B]/[C] + 1 = 1/9 +1 = 11/9